OpenStudy (anonymous):

I need help figuring out a percentage based on income. If I make 41,600 per year, and my husband makes 62,000 per year, how much do each of us pay toward a $1,000 bill for his kid to play soccer?

OpenStudy (anonymous):

i got it one second

OpenStudy (anonymous):

ok, great - thank you sooooo much!!!

OpenStudy (anonymous):

payment if received by May 27 is $75 per play this or think of it this way Garden Club donated $1000 toward the ...... bill and that interest at a rate of 6.5% per annum will be charged 60 days after ...

OpenStudy (anonymous):

around five per cent of the year for the expansion of the ..... is being taken away from us and the community must be made aware as to ... His present salary is $32500. Golf anyone? Golf Lessons will begin ...... Howler, if the bill should run into trouble

hero (hero):

41,600 + 62,000 = 103,600 Figure out what percentage of 103,600 is 41,600 Figure out what percentage of 103,600 is 62,000 Then apply those percentages to 1000

OpenStudy (anonymous):

41,600 + 62,000 = 103,601

OpenStudy (anonymous):

and then u apply the percentages to 1,000 and u got ur answer

hero (hero):

Well, first of all the person that makes the most money.....their percentage should be calculated first Husband makes most money and his percentage of income is 59.85% of their total income.. So, 59.85% of 1000 = 598.46 1000 - 598.46 = 401.54 So Husband pays $598.46 Wife pays $401.54

hero (hero):

Do you get it?

hero (hero):

Or you can just round it to the nearest ten percent and say that the husband makes 60% of the income so he pays $600 and the wife pays $400

OpenStudy (anonymous):

Ok, that makes total sense! Thank you so much. I really appreciate your help.

hero (hero):

You're welcome

hero (hero):

thanks

OpenStudy (anonymous):

perfect. Thank you sooo much!

hero (hero):

I admit, it wasn't clear at first how to figure it out. You had to somehow relate both incomes in a way that would make equivalent ratios for the 1000 dollars. If there is another way to approach this let me know.